### v1

# Compressing Scene Dynamics: A Generative Approach

Abstract
This paper proposes to learn generative priors from the motion patterns instead of video contents for generative video compression. The priors are derived from small motion dynamics in common scenes such as swinging trees in the wind and floating boat on the sea. Utilizing such compact motion priors, a novel generative scene dynamics compression framework is built to realize ultra-low bit-rate communication and high-quality reconstruction for diverse scene contents. At the encoder side, motion priors are characterized into compact representations in a dense-to-sparse manner. At the decoder side, the decoded motion priors serve as the trajectory hints for scene dynamics reconstruction via a diffusion-based flow-driven generator. The experimental results illustrate that the proposed method can achieve superior rate-distortion performance and outperform the state-of-the-art conventional video codec Versatile Video Coding (VVC) on scene dynamics sequences. The project page can be found at https://github.com/xyzysz/GNVDC.
1 Introduction
Motion is the eternal melody of the world. Efficiently characterizing motion patterns and reducing temporal redundancy is of vital importance to video coding techniques. With the recent advancements of “Artificial Intelligence Generated Content (AIGC)” techniques, generative video coding (GVC) is proposed to reduce the transmission redundancy and achieve ultra-low bit-rate coding by leveraging the statistical regularities in particular video contents such as human faces [1]. Specifically, the motion flow can be predicted from compact feature representations that are extracted from the key frame and inter frames. Then, the strong generation ability of the deep generative model can guarantee visual-pleasing reconstruction by animating the key frame with the corresponding optical flows [2]. Such a design can warrant superior Rate-Distortion (RD) performance and outperform Versatile Video Coding (VVC) with a large margin in terms of perceptual metrics [3].
Existing GVC methods usually use explicit representations with physical meaning as the precursors of motion representations. For example, Deep Animation Codec [4] adopts 2D key-points representations [5] to build generative video-conferencing codec, Face Video-to-Video Synthesis [6] further leverages 3D key-points for free-view talking-head synthesis, and an extreme generative human-oriented video codec [7] is built by compressing the articulated motion representations [8]. Meanwhile, implicit features are also explored for direct motion representations. Compact Feature Temporal Evolution [9] encapsulates temporal trajectories into $4\times 4$ matrices, and Latent Image Animation [10] extracts weighting coefficients for decomposed motion vectors. However, these methods exploit prior knowledge from video contents under the same scenario, hindering their trained models to generalize to more diverse video contents.
Recently, Generative Image Dynamics [11] illustrates the effectiveness of learning the distribution of natural motions conditioned on a given image. It can generate natural oscillatory motions across scenes using the Diffusion Model [12]. Connecting this motion prior with the prior-based generative video coding raises an intriguing question: Can we characterize specific motion pattern priors into suitable representations that are independent of video contents, so as to enable generative video coding to generalize to a wider range of scenes? To verify this idea, in this paper, we follow [11] and choose the most common motion patterns, i.e., small motion dynamics in everyday scenes, to explore the possibility of building Generative Scene Dynamics Coding framework (Dynamics-Codec). Herein, we limit “small motion dynamics” to the minor, rigid and non-articulated motion in ubiquitous daily scenes, such as oscillations and rectilinear movements of objects. It is expected that this Dynamics-Codec framework can directly extract compact feature representations from motion flows instead of video contents and realize generative reconstruction for various kinds of dynamic scenes. The difference between the video-content-prior-based and motion-pattern-prior-based generative video coding is illustrated in Figure 1.
(a) Video-content-prior-based GVC.
(b) Motion-pattern-prior-based GVC.
Figure 1: Comparison of the video-content-prior-based and motion-pattern-prior-based generative video coding.
The efforts to establish the proposed Dynamics-Codec are two-fold. First, compact motion tokens are identified from motion flows with a motion tokenizer in a dense-to-sparse manner for efficient compression. To benefit the modelling of certain motion patterns, the main characteristics of moving areas are first identified by motion sampling from the dense motions to form sparse motions. Then, the spare motions are further condensed into compact tokens. Second, a robust decoder is designed to generate reconstructions with key frame and motion information across different video contents. Herein, the Stable Video Diffusion (SVD) model [13] that is pre-trained on a wide range of data distribution is utilized to enable the generalizability to various scenes, and a pre-trained Control-Net like motion flow adaptor [14] is adopted to ensure the seamless flow-driving control in a feature-warping manner with the recovered dense motions.
The contribution of this paper can be summarized as the following,
  • We propose a novel generative scene dynamics compression framework that utilizes motion pattern priors instead of video content priors for the ultra-low bit-rate compression and high-quality reconstruction for diverse video contents.
  • We design an optical flow tokenizer that can transform dense motion between key frame and inter frame into compact representations for extremely low bit-rate compression of temporal information.
  • We develop a flow-driven decoder that can recover dense motions from the flow tokens and integrate the powerful diffusion-based generator in GVC for visual-pleasing reconstruction across various scenes.
  • The experiment results show that the proposed Dynamics-Codec can achieve superior rate-distortion performance as well as subjective quality against the state-of-the-art conventional video codec VVC.
2 The Proposed Dynamics-Codec
[…]
2.1  The Dynamics-Codec Framework
Figure 2: The detailed structure of proposed Dynamics-Codec. The networks in orange blocks are optimized from scratch, and the networks in blue block is pre-trained.
The detailed structure of proposed Dynamics-Codec is shown in Figure 2. At the encoder side, the key frame (i.e., the first frame of the video sequences) is compressed by conventional VVC codec and transmitted as image bit-stream. For inter frames, dense optical flows are first extracted between the key frame and every subsequent inter frame. Then, each dense flow is sampled by a watershed-based method [15, 16] to obtain the sparse motion and the corresponding motion mask, which are further down-sampled and vectorized to form motion tokens. To further eliminate the coding redundancy, all tokens are inter-predicted with the adjacent token, and the quantized residuals are encoded by Context Adaptive Binary Arithmetic Coding (CABAC).
At the decoder side, the reconstructed key frame is decoded from the image bit-stream, and fed into a feature extractor. Meanwhile, the motion tokens are obtained by context-based entropy decoding and feature compensation from the feature bit-stream. The dense motions are subsequently reconstructed by leveraging the internal relationship between the extracted key frame features and the decoded motion tokens. Finally, the recovered dense motions are fed into a flow-driven diffusion-based generator for denoising generation of reconstructed inter frames.
[…]
2.2  Optical Flow Tokenizer Feature representation is essential to generative video coding. Previous GVC works mainly focus on utilizing video content priors for characterizing temporal trajectory on homogeneous data such as human face or human body [2]. In this paper, we advance the idea of video-content prior to motion-pattern prior that requires direct feature extraction from optical flows. To preserve the major motion information and benefit the efficient representation, the motion tokenizer is designed as a dense-to-sparse scheme.
Specifically, dense motion is first extracted with an off-the-shelf and latest motion estimation model Unimatch [17], which utilizes a transformer-based multi-task framework to solve the unified dense correspondence matching problem. Without loss of generality, we denote the key frame and inter frame as I and P with the dimension of $3\times H\times W$, the motion extraction can be denoted as,
$\textbf{f}_{d}=\Theta(\textbf{I},\textbf{P})$ (1)
where $\Theta$ denotes motion prediction and $\textbf{f}_{d}\in\mathbb{R}^{H\times W\times 2}$ denotes the motion field in the format of coordinate grid.
To select the primary motion trajectories under the consideration of both its appearance and intensity, a topological-distance watershed map sampling [16] with motion edge detection and Non-maximum Suppression [18] is leveraged to obtain the sparse motions and the motion masks, where the latter serves to distinguish between sampled positions [15] with zero-motion and unsampled positions,
$\textbf{f}_{s},\textbf{m}_{s}=\Phi(\textbf{f}_{d})$ (2)
where $\Phi$ denotes motion sampling. $\textbf{f}_{s}\in\mathbb{R}^{H\times W\times 2}$ and $\textbf{m}_{s}\in\mathbb{R}^{H\times W\times 1}$ denote the sparse motion and the motion mask respectively. Then, a sparse motion compressor is utilized to transform sparse motions to compact motion tokens, where the motion feature is first obtained by a flow feature extractor $\textbf{e}_{f}$,
$\textbf{y}_{f}=\textbf{e}_{f}(concat[\textbf{f}_{s},\textbf{m}_{s}])$ (3)
where $\textbf{y}_{f}$ denotes the flow feature and $concat$ denotes the concatenation operation. Herein, this flow feature extractor is designed as a U-Net [19] like structure with up-sampling, down-sampling and short-cut connections. Then, two motion token vectors are derived by two vectorizers $\textbf{v}_{w}$ and $\textbf{v}_{b}$,
$\textbf{w}=\textbf{v}_{w}(\textbf{y}_{f})$ (4)
$\textbf{b}=\textbf{v}_{b}(\textbf{y}_{f})$ (5)
where $\textbf{w}\in\mathbb{R}^{N_{v}\times 1}$ and $\textbf{b}\in\mathbb{R}^{N_{v}\times 1}$. $N_{v}$ denotes the number of tokens in each vector. Here, the vectorizers are designed as cascaded convolutional layers and Generalized Divisive Normalization layers [20]. In this way, high-fidelity dense motions are transformed to very compact motion tokens, which served as high-level semantic motion hints, through the guidance of intermediate sparse motions. Such a process could exploit the compact motion pattern prior in motion flows, benefiting both high compressibility of transmitted features and decent accuracy of subsequent motion recovery.
[…]
2.3  Flow-driven Decoder In previous GVC schemes, generators are usually crafted following the structures of Generative Adversarial Network [21] or Variational Auto-Encoder [22]. The resulting GVC methods can only address single scenario application with one trained model, limiting their generalizability to more diverse contents. Recently, the Diffusion Model [12] is proposed to perform generation with the denoising process and exhibits the outstanding capability of learning complex data distribution. It can be integrated with versatile conditions for tailored generative vision tasks [23]. Motion trajectory is also leveraged as the control signal for interactive image-to-video generation [24, 25, 14, 26, 27]. However, they are not yet explored in GVC framework. In this paper, a pre-trained SVD [13] and motion-adaptor [14] are utilized to ensure the robustness of motion-driven generation and its generalizability across various contents.
Specifically, the reconstructed key frame is denoted as $\hat{\textbf{I}}$, and it is first fed into a image feature extractor $\textbf{e}_{i}$,
$\textbf{y}_{i}=\textbf{e}_{i}(\hat{\textbf{I}})$ (6)
where $\textbf{y}_{i}\in\mathbb{R}^{N_{v}\times H\times W}$ denotes the image feature. Here, the image feature extractor is designed as a U-Net like structure. The decoded motion token vector $\hat{\textbf{w}}$, $\hat{\textbf{b}}$ and the image feature are processed by a dense flow generator $\textbf{g}_{f}$,
$\hat{\textbf{m}}_{f}=\textbf{g}_{f}(\textbf{y}_{i}\cdot\hat{\textbf{w}}+\hat{\textbf{b}})$ (7)
where $\hat{\textbf{m}}_{f}\in\mathbb{R}^{H\times W\times 2}$ denotes the recovered dense motion flow and $\cdot$” denotes channel-wise multiplication. Here, the flow generator is designed as a U-Net like structure. Then, following the practice in [14], the recovered dense motion is used to warp the multi-scale features of reconstructed key frame, and the features are further fused with trainable SVD encoder features. Finally, they serve as the trajectory guidance for the pre-trained diffusion decoder. In this way, massive training data can be introduced by pre-trained models without any additional bit-rate or training cost, which can consequently guarantee the visual-pleasing reconstruction of the flow-driven decoder.
[…]
2.4  Optimization To bridge the gap between proposed motion token representations and pre-trained flow-driven diffusion-based generator, the optimization of Dynamics-Codec is aimed to align recovered dense motions with the original dense motion inputs of the motion adaptor. To achieve that, only flow tokenizer and dense motion recovery module are optimized, and the original dense motion are provided as supervision signals. Here, the original motion predictor of the motion adaptor is denoted as $\phi$, and L1 loss is used to measure the coordinate-wise error between ground truth flows and generated flows. The optimization objective can be written as,
$\mathcal{L}=||\hat{\textbf{m}}_{f}-\phi(\textbf{I},\textbf{P})||_{1}$ (8)
where $||\cdot||_{1}$ denotes the L1 norm.
3 Validations
[…]
3.1  Experimental Settings    Dataset.
Figure 3: The overview of small motion scene dynamic test set.
We collect small motion dynamic videos with various scenes and resolution of 384 $\times$ 384 from the Internet for training and evaluation. For training, total 900 videos are used to extract their ground truth dense motions between the key frame and every inter frame. For evaluation, 35 videos with diverse video contents are utilized, as shown in Figure 3, and each sequence contains 50 frame with frame rate of 25 fps.
Training Settings. We train the flow tokenizer and dense flow recovery module jointly without entropy coding. The size of image/flow features are set as 384 $\times$ 384, and the number of motion tokens in each token vector is set as 20, yielding totally 40 parameters. We implement the networks with Pytorch framework, which are then optimized by Adam optimizer with $\beta_{1}=0.5$, $\beta_{1}=0.999$ and learning rate of $10^{-4}$. We also use multi-step learning rate scheduler with $\gamma$ = 0.1 and $milestone$ = [30, 45] We train the networks on NVIDIA GeForce RTX 3090 GPU for 50 epochs with the batch size of 24.
Evaluation Settings. For Dynamics-Codec key frame compression, VVC codec with VTM 22.2 platform is used with intra mode and QP of 32, 37, 42, 47. We compare the proposed Dynamics-Codec with VVC anchor performances, which are obtained with Low-Delay-Bidirectional (LDB) configuration and QP of 37, 42, 47, 52. For evaluation metrics, we choose three perceptual measurements that are commonly used for generative contents, i.e., Learned Perceptual Image Patch Similarity (LPIPS) [28], Deep Image Structure and Texture Similarity (DISTS) [29], and Frechet Video Distance (FVD) [30]. Additionally, we use Bjøntegaard-delta-rate (BD-rate) [31] and rate-distortion (RD) curve to compare the proposed method with anchor method. To display the RD curves in an increasing manner and calculate BD-rate saving, we use “1-DISTS”, “1-LPIPS” and “5000-FVD” as the y-axis of the graphs.
[…]
(a) Rate-DISTS
(b) Rate-LPIPS
(c) Rate-FVD
Figure 4: Rate-distortion performance comparisons with VVC [32] in terms of DISTS, LPIPS and FVD for scene dynamics test set.
Table 1: RD performance comparisons on scene dynamics test set in terms of average BD-rate savings over the VVC anchor [32].
Rate-DISTS Rate-LPIPS Rate-FVD
-35.55% -23.98% -29.91%
Subjective Quality. Subjective visual quality comparisons of proposed Dynamics-Codec and VVC are shown in Figure 5. Specifically, reconstructions of test sequence 031 at 10 kbps, sequence 029 at 6 kbps and sequence 034 at 5 kbps are provided. The motion dynamics of each sequence are shown in first columns in the format of y-t space-time slices by concatenating 5-pixels-wide samples obtained from each of 50 frame. The sampled positions are indicated by red-dash lines in the first frame. The sampled dynamics present the motion patterns in sampled position and it can be seen that our method can accurately reconstruct the temporal trajectory of the objects. Then, 10 full reconstructed frames from each sequence are presented for video content comparison. It can be observed that, under ultra-low bit-rate, VVC reconstructions exist annoying blocking effects, while the Dynamics-Codec generates more visual-pleasing reconstructions with decent motion accuracy. These results demonstrate the robust generation capability of proposed motion-driven decoder for both high-quality generation and generalizability across diverse scenes.
(a) Sequence 031 at 10 kbps
(b) Sequence 029 at 6 kbps
(c) Sequence 034 at 5 kbps
Figure 5: Visual quality comparisons on test sets among VVC [32] and Dynamics-Codec (Ours) at the similar bit rate. Please refer to our project page for more demos.
4 Conclusion
In this paper, we propose to exploit motion-pattern-prior instead of video-content-prior for generative video coding. By doing so, we build a novel Dynamics-Codec for scene dynamics. In particular, the compact motion prior representations are created by identifying the primary motion components and condensing them to tokens, while the robust reconstructions among various video contents are guaranteed by powerful diffusion-based motion-driven decoder. The experiment results show that our methods can achieve more than 35% BD-rate saving against VVC as well as visual-pleasing reconstructions under ultra-low bit-rate on small motion dynamics sequences, highlighting better robustness and generalizability of the motion-prior-based generative video coding scheme.